### 12 Dec 2023 v2

@@ Abstract
Multilevel regression and poststratification (MRP) has been used extensively to adjust convenience and low-response surveys to make population and subpopulation estimates. For this method, model validation is particularly important, but recent work has suggested that simple aggregation of individual prediction errors does not give a good measure of the error of the population estimate. In this manuscript we provide a clear explanation for why this occurs, propose two scoring metrics designed specifically for this problem, and demonstrate their use in three different ways. We demonstrate that these scoring metrics correctly order models when compared to true goodness of estimate, although they do underestimate the magnitude of the score.
@@ 1 Introduction
Surveys are increasingly reliant on model-based estimates to overcome low response rates and non-probability samples. In this manuscript we focus on a model-based technique, multilevel regression and poststratification (MRP; Gelman & Little, 1997, Park et al., 2006 and Lax & Phillips, 2009b). This method has been widely used in political science (e.g., Lax & Phillips, 2009a, Ghitza & Gelman, 2013, and Wang et al., 2015) and increasingly deployed in social (e.g., Alsalti et al., 2023, and Kolczynska et al., 2020) and health sciences (e.g., Machalek et al., 2022, and Downes et al., 2018). As with all methods of survey adjustment, the method relies on assumptions of model suitability.
@@
MRP has three stages (see, e.g., Lopez-Martin et al., 2022). First, a model (originally a Bayesian multilevel model, but increasingly other regularising models, e.g., Ornstein, 2020, Bisbee, 2019, and Gao et al., 2021) is fit to the observed sample data conditional on discrete predictors (typically geographic and demographic) whose joint distribution in the population is known or can be estimated. Second, a poststratification table is constructed to represent this population joint distribution, ideally using census data or high quality official statistics. The model from the sample is then used to predict the expected value for each cell defined by $X$ in the population. Lastly, the population totals in each cell are used to weight the cell estimates to obtain an estimate for the population (or subpopulation). The three steps can be remembered as model-predict-aggregate.
@@
The goodness of the population estimate will be heavily reliant on the goodness of the model predictions. This means that it is essential to consider how to fairly validate the model. Given the emphasis on prediction in this method, previous work investigated the efficacy of the sum of leave-one-out(LOO) expected log predictive density (ELPD) in estimating true error (Kuh et al., 2022). This is a commonly used metric for model goodness in the Bayesian literature, and is easily implemented through the LOO package (Vehtari et al., 2021). Kuh et al. (2022) also considered a weighted alternative to this metric (initially proposed by Lumley & Scott, 2015) to account for the non-random sample.
Little & Vartivarian (2005) demonstrate that the required adjustment variables to produce an unbiased estimate are those that are related to both the probability of inclusion and the probability of the outcome (which Kuh et al., 2022 refer to as bias variables). However, neither standard sum of LOO ELPD nor the weighted alternative correctly distinguish between these bias variables and precision variables (variables with are strongly related to the model outcome but not sample inclusion). This is a particular challenge for MRP, as each variable in the model needs to be found or estimated in the population in order to create a poststratification matrix. Kuh et al. (2022) also note additional challenges in this space, such as the difference in preferred model for different subpopulations.
@@ 1.1 Individual scores and aggregate scores
We argue that the underlying difficulty described in Kuh et al. (2022) is that the best model for an average over individuals may not correspond to the best model for an aggregate estimate. To understand why, consider a simple example of a sample of size 2, where these individuals have a true outcome value 0. We will say $y_{\mathrm{true}}=\{0,0\}$. The mean of $y_{true}$ is $E(y_{\mathrm{true}})=\frac{0+0}{2}=0$. Now we have two potential models to estimate $y_{\mathrm{true}}$. Each produces an estimate (column 2, Table 1), from which we can calculate the squared error and the mean squared error over both individuals. We can also use these estimates to create an estimate of the mean, and from this calculate the squared error of the mean. As we can see, while model 1 is preferred by the mean squared individual error, model 2 is far preferred at estimating the mean.
@@
This simple example demonstrates what we believe to be one of the underlying challenges of validating multilevel regression and poststratification (MRP) models. An MRP based estimator is by definition an estimator that uses cellwise estimates to compose a population or subpopulation estimates. Previously we had investigated using sum of individual goodness to estimate the goodness of the MRP estimate, which is not effective (Kuh et al., 2022).
@@ 1.2 Notation and key terms
Consider a finite population $P$ with $N$ individuals, where interest lies in a binary variable $Y$. Whilst some authors have used a continuous outcome variable (Liu et al., 2023), the prevalent applications currently focus on estimating probabilities. Assume also a set of $k$ predictors whose values are known in the population and are represented by a $N\times k$ matrix, $\mathbf{X}$. Assume the predictors are discrete, with the $k^{th}$ variable, $X^{(k)}$, having $l^{(k)}$ levels. This means the distribution of $X$ in the population can be represented by the number of items within each of the $J=\prod_{k}l^{(k)}$ possible configurations of predictors.
@@
$\operatorname{E}_{\mathrm{P}}(Y)=\frac{\sum_{i=1}^{N}Y_{i}}{N}.=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1)}{N},$ (1)
@@
Next take a sample of size $n$. The observations in this sample are assumed to be random conditional on the predictors, $X$. The probability of inclusion in the sample for every $j^{th}$ cell is denoted $\pi_{j}$.
Following traditional sampling notation, we use lower case for the sample and upper case for the population, and where a matrix is needed, it will be presented in boldface. Our goal is to estimate the population mean by first estimating the probability of the outcome in a each cell in the population using the sample. We denote this quantity for the $j^{t}h$ cell as $\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\theta)$. We then use this to estimate the population mean, $E_{P}(Y|y,\textbf{x},\theta)$. We do this by modifying (1) into
$\operatorname{E}_{\mathrm{P}}(Y|y,\textbf{x},\theta)=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\theta)}{\sum_{j=1}^{J}N_{j}}.$ (2)
Our goal is to identify the best model $M$ and estimate the parameters $\theta$, denoted $\hat{\theta}$, using the sample. We can then use this to estimate $\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})$, which can then be used in (2).
$\operatorname{E}_{\mathrm{P}}(Y|y,\textbf{x},\hat{\theta})=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})}{\sum_{j=1}^{J}N_{j}}.$ (3)
In this paper we discuss aggregation at a range of different levels. For clarity we describe all terms with a simple example. Consider a survey of voting intention. The smallest unit is the individual, the person who is being surveyed. These individuals can be aggregated in different ways. For example, we could aggregate all individuals with a specific set of demographic qualities (e.g., aged 18–25, female, college educated). We call this a cell if these three demographics are the full set of $X$ variables. We could also aggregate at a particular level (e.g., 18-25 year) of a variable (age). This is a special case of a subpopulation (also known as a small area), which is where we estimate part of the population. In the limit the subpopulation is the full population (e.g., all US voters).
@@ 1.3 Contributions
  • Propose an alternative scoring metric when estimating the probability of a binary outcome in the population (Section 2)
@@
  • The leave-one-cell-out approach requires all cells to be observed in the sample. We also propose an approximate leave-one-cell-out reference approach for use when not all cells are observed in the sample (Section 5), and a combined validation approach that incorporates leave-one-cell-out cross validation and leave-one-cell-out reference validation (Section 5).
Together we intend for these contributions to provide a convincing argument that the proposed scoring metric is more appropriate for scoring models for MRP. The structure of this paper is as follows. In Section 2 we describe our scoring rule adjustment. In Section 3 we apply our scoring rules in a leave-one-cell-out cross validation scheme. In Section 4 we apply our scoring rules to score subpopulation estimates using approximate leave-one-cell-out cross validation. In Section 5 we discuss the use of reference validation scores.
2 Aggregate scoring from cellwise error
In this section we demonstrate that by decomposing a score of the MRP population mean into its cellwise components, we can produce an estimate of the population mean score using cellwise goodness. We complete this for a squared error score ($\operatorname{Error}^{2}$) for the population mean and the continuous ranked probability score (CRPS). This decomposition may be possible for other scoring rules.
@@ 2.1 Squared Error
First consider the squared error for the population mean,
$\operatorname{Error}_{\mathrm{MRP}}^{2}=(E_{P}(Y|\theta,y,\textbf{x})-E_{P}(Y))^{2}.$ (4)
@@
$\operatorname{Error}_{\mathrm{MRP}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})}{\sum_{j=1}^{J}N_{j}}-\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1)}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2},$ (5)
@@
$\operatorname{Error}_{\mathrm{MRP}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})\,-\,\mbox{Pr}(Y_{i\in j}=1))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2}.$ (6)
The sum is on the inside of the squared term, which demonstrates why simply summing the squared error of the individuals or cells (as is often done when scoring models; see Gelman et al., 2014) would not provide an acceptable score for the population mean. For clarity, the mean of the squared error of the cells, adjusted for sample representation would be
$\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\hat{\theta})-\mbox{Pr}(Y_{i\in j}=1))^{2}}{\sum_{j=1}^{J}N_{j}}.$ (7)
@@ 2.2 Continuous ranked probability scores
where $F$ is the CDF of the predictive distribution and $x$ is the true value (Gneiting & Raftery, 2007). One benefit of this score is that it can be implemented through approximate draws of this distribution (Equation (21) of Gneiting & Raftery, 2007). To harmonise notation, we denote $\phi$ as the posterior estimate for the population expectation.
$\operatorname{CRPS}(\phi,E_{P}(Y))=\frac{1}{2}E(\lvert\phi-\phi^{\prime}\rvert)-E(\lvert\phi-E_{P}(Y)\rvert).$ (9)
In our setting we focus on cases where we use posterior draws to estimate expectation. Denoting the $b^{th}$ posterior draw as $\mathbf{\phi}^{b}$ and $\mathbf{\phi}^{\prime}$ as a permutation of the posterior draws of $\mathbf{\phi}$, we can write this as
@@
The $\phi$ represents our MRP estimate, and $E_{P}(Y)$ the population mean truth, so we can rewrite this as
$\!\!\!\!\!\!\!\!\operatorname{CRPS}_{\mathrm{MRP}}(Y,y,\hat{\theta},\mathbf{x})=$
$\frac{1}{B}\sum_{b}^{B}\Bigg{(}\frac{1}{2}\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}(\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{b},\textbf{x})*N_{j})}{\sum_{j=1}^{J}N_{j}}\,-\,\frac{\sum_{j=1}^{J}(\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{\prime b},\textbf{x})*N_{j})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\,-$
$\ \ \ \ \ \ \ \ \Bigg{\lvert}\frac{\sum_{j=1}^{J}(\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{b},\textbf{x})*N_{j})}{\sum_{j=1}^{J}N_{j}}\,-\,\frac{\sum_{j=1}^{J}(\mbox{Pr}(Y_{i\in j}=1)*N_{j})}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$
@@
$\!\!\!\!\operatorname{CRPS}_{\mathrm{MRP}}(Y,y,\hat{\theta},\mathbf{x})=\frac{1}{B}\sum_{b}^{B}\Bigg{(}\frac{1}{2}\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{\prime b},\textbf{x})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\,-\\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}=1|y,\hat{\theta}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j}=1)\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (10)
As with squared error, the absolute value is outside of the summation, rather than inside as we would expect for a usual sum of the cellwise scores:
$\begin{split}\operatorname{CRPS}(Y,y,\hat{\theta},\mathbf{x})=\frac{1}{N}\sum_{j=1}^{J}N_{j}\Bigg{(}\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Big{\lvert}\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{\prime b})\Big{\rvert}\,-\,\\ \frac{1}{B}\sum_{b}^{B}\Big{\lvert}\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{b})-\mbox{Pr}(Y_{j})\Big{\rvert}\Bigg{)}.\end{split}$ (11)
2.3 Demonstration of equivalence
Here we demonstrate the equivalency of the cellwise decomposition to the original population score. To achieve this, we directly calculate the squared error and CRPS for the MRP estimate at the population level and then also calculate this using the cellwise estimates.
For simplicity, we borrow the simulation scenario in Example 1 of Kuh et al. (2022), but add an additional constraint that all cells in the population must be observed at least once in the sample (Kuh et al., 2022 had a less restrictive constraint that at least one observation every level needed to be observed). In this example, four independent, normally distributed ($\mbox{normal}(\mu=0,\sigma=2)$) variables are created. The probability of the outcome ($\mbox{Pr}(Y_{i}=1)$) and the probability of being included in the sample ($\pi_{i}$) were created using the following relationships:

### 12 Apr 2024 v3

@@ Abstract
Generalization to new samples is a fundamental rationale for statistical modeling. For this purpose, model validation is particularly important, but recent work in survey inference has suggested that simple aggregation of individual prediction scores does not give a good measure of the score for population aggregate estimates. In this manuscript we explain why this occurs, propose two scoring metrics designed specifically for this problem, and demonstrate their use in three different ways. We show that these scoring metrics correctly order models when compared to the true score, although they do underestimate the magnitude of the score. We demonstrate with a problem in survey research, where multilevel regression and poststratification (MRP) has been used extensively to adjust convenience and low-response surveys to make population and subpopulation estimates.
@@ 1 Introduction
A central challenge of statistics is to generalizing from observed data to new settings. To do this, sample surveys increasingly rely on model-based estimates to overcome low response rates and non-probability samples. These model-based estimates use a prediction framework, but the purpose is not to predict each unobserved individual accurately. Instead the purpose is to estimate accurate population summaries from individual predictions. Although these two aims might seem similar, recent work suggests that the best model for the individual predictions might not be the best model for population summaries (Kuh et al., 2023). This leads to challenges in assessing both model adequacy and model selection (Gelfand et al., 1992).
@@
In this manuscript we focus on a specific model-based technique, multilevel regression and poststratification (MRP; Gelman & Little, 1997; Park et al., 2006), which has been widely used in political science (e.g., Lax & Phillips, 2009b, a; Ghitza & Gelman, 2013; Wang et al., 2015) and increasingly deployed in social (e.g., Alsalti et al., 2023; Kolczynska et al., 2024) and health sciences (e.g., Machalek et al., 2022; Downes et al., 2018). As with all methods of survey adjustment, the method relies on assumptions of model suitability.
@@
MRP has three stages (see, e.g., Lopez-Martin et al., 2022). First, a model (originally a Bayesian multilevel model, but increasingly other regularising models, e.g., Ornstein, 2020; Bisbee, 2019; Gao et al., 2021) is fit to the observed sample data conditional on discrete predictors (typically geographic and demographic) whose joint distribution in the population is known or can be estimated. Second, a poststratification table is constructed to represent this population joint distribution, ideally using counts derived from census data or high quality official statistics. The model from the sample is then used to predict the expected value for each cell. Third, the population totals in each cell are used to weight the cell estimates to obtain an estimate for the population (or subpopulation). The three steps can be remembered as model-predict-aggregate.
@@
The population estimate relies on the model prediction, and so it is essential to consider how to fairly validate the model. Given the emphasis on prediction in this method, previous work has investigated the sum of leave-one-out (LOO) expected-log-predictive-density (ELPD) (Kuh et al., 2023). This is a commonly used metric for model goodness in the Bayesian literature, and is easily implemented through the LOO package (Vehtari et al., 2021). Kuh et al. (2023) also consider a weighted alternative to this metric (initially proposed by Lumley & Scott, 2015) to account for non-random sampling.
Kuh et al. (2023) use results from Little & Vartivarian (2005) to demonstrate that neither standard sum of LOO-ELPD nor the weighted alternative correctly distinguish between a model that correctly produces an approximately unbiased estimate of the mean and one that does not.
@@ 1.1 Individual scores and aggregate scores
We argue that the underlying difficulty described in Kuh et al. (2023) is that the best model for predicting individuals may not correspond to the best model for an aggregate estimate. To understand why, consider a simple example of a sample of size 2, where these individuals have a true outcome value 0. We will say $y_{\mathrm{true}}=\{0,0\}$. The mean of $y_{true}$ is $E(y_{\mathrm{true}})=\frac{0+0}{2}=0$. Now we have two potential models to estimate $y_{\mathrm{true}}$. Each produces an estimate (column 2, Table 1), from which we can calculate the squared error and the mean squared error over both individuals. We can also use these estimates to create an estimate of the mean, and from this calculate the squared error of the mean. As we can see, while model 1 is preferred by the mean squared individual error, model 2 is far preferred at estimating the mean.
@@
This simple example demonstrates what we believe to be one of the underlying challenges of validating multilevel regression and poststratification (MRP) models. An MRP based estimator is by definition an estimator that uses cellwise estimates to compose a population or subpopulation estimates. Previous investigations used sum of individual goodness to estimate the goodness of the MRP estimate, which was not effective (Kuh et al., 2023).
@@ 1.2 Notation and key terms
Consider a finite population $P$ with $N$ individuals, where interest lies in a binary variable $Y$ [^4] Whilst some authors have used a continuous outcome variable (Liu et al., 2023), the prevalent applications currently focus on estimating probabilities.. Assume a set of $k$ predictors whose values are known in the population and are represented by a $N\times k$ matrix, $\mathbf{X}$. Assume the predictors are discrete, with the $k^{th}$ variable, $X^{(k)}$, having $l^{(k)}$ levels. This means the distribution of $X$ in the population can be represented by the number of items within each of the $J=\prod_{k}l^{(k)}$ possible configurations of predictors.
@@
$\operatorname{E}_{\mathrm{P}}(Y)=\frac{\sum_{i=1}^{N}Y_{i}}{N}=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}=1)}{N},$ (1)
@@
Next take a sample of size $n$. The probability of inclusion in the sample for every $j^{th}$ cell is assumed to be the same and denoted $\pi_{j}$.
Following traditional sampling notation, we use lower case for the sample and upper case for the population, and where a matrix is needed, it will be presented in boldface or evident from context. Our first step is to estimate the probability of the outcome in the $j^{th}$ cell in the population using the sample, denoted $\mbox{Pr}(Y_{i\in j}|y,\theta)$ as shorthand for $\mbox{Pr}(Y_{i\in j}=1|y,\textbf{x},\theta)$. We then use this to estimate the population mean, $E_{P}(Y|y,\theta)$. We do this by modifying (1) into
$\operatorname{E}_{\mathrm{P}}(Y|y,\theta)=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}|y,\theta)}{\sum_{j=1}^{J}N_{j}}.$ (2)
Our goal is to identify the best model $M$ and estimate the parameters $\theta$, denoted $\hat{\theta}$, using the sample. We can then use this to estimate $\mbox{Pr}(Y_{i\in j}|y,\hat{\theta})$, which can then be used in (2).
$\operatorname{E}_{\mathrm{P}}(Y|y,\hat{\theta})=\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta})}{\sum_{j=1}^{J}N_{j}}.$ (3)
In this paper we discuss aggregation at a range of different levels. For clarity we describe all terms with a simple example. Consider a survey of voting intention. The smallest unit is the individual, the person who is being surveyed. These individuals can be aggregated in different ways. For example, we could aggregate all individuals with a specific set of demographic qualities (e.g., aged 18–25, female, college educated). We call this a cell if these three demographics are the full set of $X$ variables. We could also aggregate at a particular level (e.g., 18-25 year) of a variable (age), which we call a subpopulation (also known as a small area). In the limit the subpopulation is the full population (e.g., all US voters).
@@ 1.3 Contributions
  • Propose two alternative scoring metrics for estimating the probability of a binary outcome in the population (Section 2)
@@
  • The leave-one-cell-out approach requires all cells to be observed in the sample. We also propose an approximate leave-one-cell-out reference approach for use when not all cells are observed in the sample (Section 5).
Together we intend for these contributions to provide a convincing argument that the proposed scoring metric is more appropriate for scoring models for MRP.
2 Aggregate scoring from cellwise components
In this section we demonstrate that by decomposing a score of the MRP population mean into its cellwise components, we can produce an estimate of the population mean score using cellwise goodness. We complete this for a squared error score ($\operatorname{Error}^{2}$) for the population mean and the continuous ranked probability score (CRPS).
@@ 2.1 Squared Error
First consider the square of the error for the population mean,
$\operatorname{Error}_{\mathrm{MRP}}=E_{P}(Y|\theta,y)-E_{P}(Y).$
@@
$\operatorname{Error}_{\mathrm{MRP}}=\ \frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta})}{\sum_{j=1}^{J}N_{j}}-\frac{\sum_{j=1}^{J}N_{j}\mbox{Pr}(Y_{i\in j})}{\sum_{j=1}^{J}N_{j}}\,$
@@
$\operatorname{Error}_{\mathrm{MRP}}=\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}|y,\hat{\theta})\,-\,\mbox{Pr}(Y_{i\in j}))}{\sum_{j=1}^{J}N_{j}}.$ (4)
This sum itself needs to be squared to get the squared error for the population mean
$\operatorname{Error}_{\mathrm{MRP}}^{2}=\Bigg{(}\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}|y,\hat{\theta})\,-\,\mbox{Pr}(Y_{i\in j}))}{\sum_{j=1}^{J}N_{j}}\Bigg{)}^{2},$
which is why simply summing the squared error of the individuals or cells (as is often done when scoring models; see Gelman et al., 2014) would not provide an acceptable score for the population mean. For clarity, the mean of the squared error of the cells, adjusted for sample representation would be
$\frac{\sum_{j=1}^{J}N_{j}(\mbox{Pr}(Y_{i\in j}|y,\hat{\theta})-\mbox{Pr}(Y_{i\in j}))^{2}}{\sum_{j=1}^{J}N_{j}}.$
For notation simplicity we write the error instead of squared error for formulae throughout this manuscript, but figures show the squared error.
@@ 2.2 Continuous ranked probability scores
where $F$ is the cumulative distirbution function of the predictive distribution and $x$ is the true value. One benefit of this score is that it can be implemented through approximate draws of this distribution (Equation (21) of Gneiting & Raftery, 2007). To harmonise notation, we denote $\phi$ as the posterior estimate for the population expectation and $\mathbf{\phi}^{\prime}$ as a permutation of the posterior draws of $\mathbf{\phi}$,
$\operatorname{CRPS}(\phi,E_{P}(Y))=\frac{1}{2}E(\lvert\phi-\phi^{\prime}\rvert)-E(\lvert\phi-E_{P}(Y)\rvert).$
In our setting we focus on cases where we use posterior draws to estimate expectation. Denoting the $b^{th}$ posterior draw as $\mathbf{\phi}^{b}$, we can write this as
@@
The $\phi$ represents our MRP estimate, and $E_{P}(Y)$ the population mean truth, so we can rewrite as
$\!\!\!\!\!\!\!\!\operatorname{CRPS}_{\mathrm{MRP}}(Y,y,\hat{\theta})=$
$\frac{1}{B}\sum_{b}^{B}\Bigg{(}\frac{1}{2}\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}(N_{j}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b}))}{\sum_{j=1}^{J}N_{j}}\,-\,\frac{\sum_{j=1}^{J}(N_{j}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{\prime b}))}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\,-$
$\ \ \ \ \ \ \ \ \Bigg{\lvert}\frac{\sum_{j=1}^{J}(N_{j}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b}))}{\sum_{j=1}^{J}N_{j}}\,-\,\frac{\sum_{j=1}^{J}(N_{j}\mbox{Pr}(Y_{i\in j}))}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$
@@
$\!\!\!\!\operatorname{CRPS}_{\mathrm{MRP}}(Y,y,\hat{\theta})=\frac{1}{B}\sum_{b}^{B}\Bigg{(}\frac{1}{2}\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{\prime b})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\,-\\ \Bigg{\lvert}\frac{\sum_{j=1}^{J}N_{j}\Big{(}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j})\Big{)}}{\sum_{j=1}^{J}N_{j}}\Bigg{\rvert}\Bigg{)}.$ (5)
As with squared error, the absolute value is outside of the summation over cells, rather than inside as we would expect for a usual sum of the cellwise scores:
$\begin{split}\operatorname{CRPS}(Y,y,\hat{\theta})=\frac{1}{N}\sum_{j=1}^{J}N_{j}\Bigg{(}\frac{1}{2}\frac{1}{B}\sum_{b}^{B}\Big{\lvert}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{\prime b})\Big{\rvert}\,-\,\\ \frac{1}{B}\sum_{b}^{B}\Big{\lvert}\mbox{Pr}(Y_{i\in j}|y,\hat{\theta}^{b})-\mbox{Pr}(Y_{i\in j})\Big{\rvert}\Bigg{)}.\end{split}$ (6)
3 Scoring rule applied with cross validation
We aim to score our model without the population truth through the observed values. We first consider a scenario where at least one individual in every cell is observed. We continue this assumption in Section 4 where we consider subpopulation estimates. In Section 5.2 we relax this assumption with a reference validation approach.
3.1 Simulation Design
For simplicity, we borrow the simulation scenario in Example 1 of Kuh et al. (2023), but add an additional constraint that all cells in the population must be observed at least once in the sample. Kuh et al. (2023) had a less restrictive constraint that at least one observation in every level needed to be observed. To create our simulated data, first four independent $\mbox{normal}(0,2)$ variables are sampled. The probability density of the outcome ($\mbox{Pr}(Y_{i\in j=1})$) and the probability of being included in the sample ($\pi_{i\in j}$) were created as follows: